The angles of a triangle are in A.P series.if the smallest ar medium angles are converted into degrees, their ratio is 3:5 . Evaluate the value of angles in radian

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the angles be a-d,a,a+d
Sum of all Angles of Triangle = 180
3a = 180
a = 60 degrees
Smallest angle = a-d
Medium Angle= a
Given, (a-d)/a=3/5
1-(d/a) = 3/5
(d/a) = 2/5
d= (2/5) x 60 = 24 degrees
Hence, The angles in degrees will be 36 , 60 , 84.
Converting into radians
The angles will be (pi/5) , ( pi/3) , ( 7 pi/15)

Thanks